LootAtTheTime:

Could someone explain how to find a driectrix of a parabola? X=\frac{1}{8}Y^2

LootAtTheTime:

@Vocaloid

Vocaloid:

first multiply both sides by 8 to convert to standard form 8x = y^2 then compare to the standard form: 4px = y^2 so if we put 8x = y^2 and 4px = y^2 side by side we see that 8x = 4px making p = 2 the directrix = -p = -2 = your sol'n

Vocaloid:



Vocaloid:

as an addendum this method only works if the vertex is right at (0,0) like it is in our problem

LootAtTheTime:

I see that when I look at it when I graph it so this method would only work if the vertex is at zero?

Vocaloid:

yeah if it's not at the vertex, you will perform a similar calculation (to find p) but there are a few more steps

Vocaloid:

(h,k) are the coordinates of the vertices (which you will need to find from the equation), then the focus is (h + p,k) for sideways parabolas and (h, k+p) for up/down parabolas

LootAtTheTime:

So when you say compare to the standard form what are you meaning?

LootAtTheTime:

Like subtraction it from the original or dividing it?

Vocaloid:

like, when we put these two expressions side by side 8x = y^2 and 4px = y^2 we can see that 8x must be equal to 4px

Vocaloid:

"compare to standard form" just means take the equation and put it side by side with the standard form equation 4px = y^2 to see which parts correspond to each other

Vocaloid:

the "4px" and the "8x" parts correspond, so we can set them equal to each other which is how we solve for p

LootAtTheTime:

im just so lost to where you got 4x I cant wrap my head around it. Im sorry I have trouble understanding these things sometimes.

Vocaloid:

"4px = y^2" is just the standard form equation for a parabola, you just have to know/memorize this the equation being discussed in the problem is 8x = y^2 since both equations equal y^2 we know that 8x and 4px are equal hope that makes a little more sense

LootAtTheTime:

Oh, ok I totally understand now. I kind of figured that was the case, and thanks this has been super helpful so far, and im huessing you just get rid of the y^2 then correct?

Vocaloid:

yeah sort of, it's like this: if A = B and C = B then A and C are equal since they're both equal to be same logic with our problem, 4px and 8x are both equal to y^2 so 4px = 8x

Vocaloid:

*equal to B

LootAtTheTime:

Thank you so much :D this has really helped a ton!